OpenStudy (blackstreet23):

Shakina and Juliette set the car's initial velocity to zero and set the acceleration to +1.2 m/s2, then clicked "start." Answer the following questions. What is the car's displacement between 0 and 10.0 s? m What was the total distance traveled by the car during this time interval? m

OpenStudy (blackstreet23):

@PlasmaFuzer @Shalante @freckles

OpenStudy (blackstreet23):

@geerky42

OpenStudy (blackstreet23):

they start at X = -20

OpenStudy (anonymous):

Hello again :D

OpenStudy (blackstreet23):

ohh sorry this one was done with a simulator of the online tool of my school

OpenStudy (blackstreet23):

there is the last one. I guess...

OpenStudy (anonymous):

Ok so are these the givens: Acceleration is constant given by: \[ a(t)=a=1.2 \frac{m}{s^2} \] Iinitial velocity: \[ v_0=0\frac{m}{s}\] and Initial position: \[ x_{0}=-20 m \]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

correct?

OpenStudy (blackstreet23):

could you please help me with the last question i posted

OpenStudy (blackstreet23):

i mentioned you in it

OpenStudy (anonymous):

Oh not this one?

OpenStudy (blackstreet23):

well later. But i submited this one already

OpenStudy (blackstreet23):

I am trying to finish this quickly so i can study for my Calc class too haha

OpenStudy (anonymous):

Oh ok... it is fairly simple if you know how to do the integrations... ok Ill go to the other one.... see you there

OpenStudy (blackstreet23):

I mean this one

OpenStudy (anonymous):

Hello again

OpenStudy (anonymous):

When u get back let me know if those are the correct initial parameters

OpenStudy (blackstreet23):

Hey. I am so sleepy. Thanks for everything. I will go to sleep now. Again you are awesome ! I really appreciate all your help!

OpenStudy (anonymous):

oh ok sorry i couldnt help here :( ill try and post some hints to point u in the right direction when u get back

OpenStudy (anonymous):

Hello by any chance are you on?

OpenStudy (blackstreet23):

Yes I am on. BTW do you know calculu?s

OpenStudy (anonymous):

yes hold on one moment sorry :/

OpenStudy (anonymous):

Sorry I have a couple things going on.... ok yes what is your question? Besides the obvious one (this problem)

OpenStudy (anonymous):

Yes I know calculus quite well

OpenStudy (anonymous):

Let me know when you can sit down for a minute because I want to try and talk you through this in one go

OpenStudy (blackstreet23):

Ohh I mentioned you in my calculus problem. Could you take a look at it?

OpenStudy (anonymous):

Oh wait it isnt this problem?

OpenStudy (anonymous):

im in the other problem